# Gravitational wave probes of particle dark matter: a review

### v4

@@ II.6.1 Minimally coupled dark matter
A commonly studied example of a minimally coupled scalar DM field $\phi$ is described by the Lagrangian Khmelnitsky and Rubakov (2014):
@@
Scalar DM can cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky and Rubakov (2014), i.e. introduce stochastic fluctuations of the metric Kim and Mitridate (2024); Kim (2023). Such a difference in the arrival times can be written as Khmelnitsky and Rubakov (2014); Porayko et al. (2018):
@@ II.6.2 Non-minimally coupled (conformal) dark matter
Another form of DM that couples only to gravity would be conformal DM. This kind of DM typically involve scalar fields that couple to the trace of the energy-momentum tensor. These models are closely related to the dilaton-type scalar DM discussed in Section II.3, but with specific couplings motivated by scalar–tensor theories of gravity, such as Fierz–Jordan–Brans–Dicke (FJBD) and Damour–Esposito–Farese (DEF) theories Fierz and Pauli (1939); Fierz (1956); Dicke (1962); Jordan (1959); Damour and Esposito-Farese (1992, 1993).
In such models, DM couples universally to the gravitational sector through a conformal rescaling of the metric. This coupling can induce a gravity-mediated force between neutron stars and the ultralight dark matter field, altering the moment of inertia of rotating pulsars – a manifestation of the Nordtvedt effect Nordtvedt (1968). These variations lead to deterministic changes in pulsar ToA and constitute a violation of the equivalence principle. This kind of conformal scalar DM has been extensively studied and constrained using binary pulsar timing measurements Damour and Taylor (1992); Kramer et al. (2006), which will be discussed inSection III.3.
@@
where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz and Barausse (2024) and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, $\theta(\vec{x})$ is a random phase, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered (see Section III.3 for a description of these terms).
@@ III.1 Methods
As explained in Section II.1, the expected signal frequency is fixed by the DM mass, but has some stochastic variations of $\mathcal{O}(10^{-6})$ Centers et al. (2021). Because the interferometers always exist in the DM field, the signal is always present. Thus, we can describe the expected signal as: quasi-monochromatic and quasi-infinite duration. Essentially, we are looking for a resonance at a particular frequency fixed by the DM mass.
@@ III.1.1 Cross-correlation
Conceptually, cross-correlation Pierce et al. (2018); Guo et al. (2019) requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin. The cross power is then divided by the auto-power (the power spectral density) of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency. It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio (SNR).
Mathematically, the cross-correlated signal strength for detector pair $IJ$ is Pierce et al. (2018); Guo et al. (2019):
@@
In addition to the signal strength, we have to obtain an estimate of the variance of the noise data. This can be computed from the individual power spectral densities in each FFT as Guo et al. (2019):
@@
Depending on the type of ultralight dark matter being searched for, the detectability of a signal in cross-correlation searches between interferometers depends on the spin of the DM: spin-0 (scalar), spin-1 (vector), or spin-2 (tensor) Manita et al. (2024). The SNR for a given detector pair is determined by the DM energy density fraction in that spin state, the strength of its couplings to standard-model particles, the observation time, and the relative orientation, location and geometry of the detectors. The latter is characterized by an “overlap reduction function” (ORF), which determines the sensitivity loss when applying cross-correlation to non-collocated, non-aligned detector pairs. For these models of DM, two kinds of ORFs arise: one from the spatial displacement of mirrors, and another from the finite-light time travel effects of the signal on the mirrors, a previously explained in Section II.4 Morisaki et al. (2021). In particular, cross-correlation is not nearly as sensitive to the finite-light travel time effect as to the spatial effect when it was applied to search for spin-1 DM, which will be shown in Section III.2.
Larger couplings, longer effective observation times, and favorable overlap between detectors increase the SNR and thus the probability of detecting a signal. Conversely, weaker couplings, shorter observations, or misaligned detectors reduce the SNR, making a signal more difficult to distinguish from instrumental noise. After an SNR is computed via Eq. 68, it can be translated directly from SNR to coupling strengths using Eq. 120-122 in Manita et al. (2024)).
@@
$h_{0,j}=\sqrt{\frac{2\,\mathrm{SNR}\,P_{I,j}\,P_{J,j}}{\gamma^{2}\,T_{\text{obs}}\,T_{\text{FFT}}}},$ (68)
where $\gamma$ is the ORF that depends on the way that different kinds of DM couple to the interferometers Manita et al. (2024). By inserting the SNR returned from a search (Eq. 67), we can obtain an estimate of the minimum detectable strain amplitude for a potential DM signal, and then map those amplitude values to constraints on DM couplings in spin-0, spin-1 or spin-2 using the strain amplitude equations derived in Section II, analogously to translating the SNR directly to coupling strengths in Manita et al. (2024).
III.1.2 Excess power method
Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass. Thus, another method Miller et al. (2021a) was developed that can vary $T_{\text{FFT}}$ as a function of frequency, and that is more robust against noise disturbances. It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their equalized powers are (1) above a given threshold $\theta_{\rm thr}=2.5$ and (2) local maxima. The equalized power is computed at each time and frequency bin by calculating the ratio $R_{ij}$ of the square modulus of the FFT with a running median estimation of the power spectral density, as in Section III.1.1:
@@ III.1.2 Excess power method
where $n$ is the number of peaks at a given frequency, and $\mu$ and $\sigma$ are the mean and standard deviation of the number of peaks in the histogram. The $CR$, like the SNR, follows a normal distribution. Using the $CR$ and assuming Gaussian noise, the minimum strain amplitude of a sinusoidal signal can be derived that would, in a frequentest interpretation, produce a detectable signal in a fraction $\geq\Gamma$ of a repeated number of experiments Astone et al. (2014); Miller et al. (2021a):
@@
The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations, and are valid for ground-based L-shaped GW interferometers (for triangular shapes, the averages will be slightly different). The $\frac{\sqrt{2}}{3}$ factor comes from an average over the DM Maxwell-Boltzmann distributed velocities as discussed in Section II.4, while the $\frac{2\pi}{2.4308}$ factor comes from convolving a sinusoid with a rectangular window function when computing the sensitivity of GW interferometers to a sinusoidal signal Astone et al. (2014).
@@ III.1.3 Logarithmic power spectral density (LPSD) method
To maximize sensitivity to ultralight dark-matter signals, it is important to let $T_{\text{FFT}}$ match $T_{\text{coh}}$ as the DM mass varies. The logarithmic power spectral density (LPSD) method Tröbs and Heinzel (2006); Vermeulen et al. (2021); Göttel et al. (2024) implements this idea by constructing spectra that are logarithmically spaced in frequency, in contrast to the excess power method in Section III.1.2 that use linearly spaced spectra every 1 Hz. While this approach is computationally intensive due to the need for a separate implementation from the standard discrete FFT, it captures subtle features in the data that could otherwise be missed. In particular, since ultralight dark matter induces fractional frequency variations on the order of $\mathcal{O}(10^{-6})$, a logarithmic spectral estimate is well-suited to resolving such tiny modulations.
@@
By applying Eq. 74 to the logarithmically estimated spectra, one can extract an estimate of the strain $h(\omega)$ from the data. This estimate can then be mapped to constraints on the coupling constants $\Lambda_{\gamma}$ and $\Lambda_{e}$, providing a direct connection between the interferometer measurement and the underlying ultralight dark matter model Göttel et al. (2024); Cowan et al. (2011). In this way, the LPSD method offers a way to detect logarithmically-spaced frequency modulations that might be missed by linear spectral analyses.
@@ III.1.4 Stochastic summing method
The stochastic summing method Nakatsuka et al. (2023) exploits the fact that when $T_{\text{FFT}}>T_{\text{coh}}$, the signal power from ultralight dark matter is spread stochastically over a range of frequencies rather than concentrated at a single frequency. Instead of searching for a narrow peak, this method sums the power across all frequency bins where the signal is expected, producing a detection statistic that captures the total available signal power.
@@
The sensitivity of the stochastic summing method has been studied for both axions and dark photons Nakatsuka et al. (2023), using a combination of analytical estimates and numerical simulations. In the absence of a signal and assuming Gaussian noise, a detection threshold can be established based on a chosen false-alarm probability. This threshold can then be used to determine, at a given confidence level, the minimum detectable signal amplitude, similar to Eq. 71 in Section III.1.2.
The exact sensitivity depends on the spectral shape of the signal, which in turn is influenced by the relative motion of the DM with respect to the interferometer. Conservative and optimistic assumptions about the direction of the DM velocity lead to $\mathcal{O}(1)$ variations in the sensitivity, which are accounted for in this method Nakatsuka et al. (2023).
Once the minimum detectable strain amplitude is determined, it can be translated into upper limits on the relevant coupling constants for each type of dark matter interaction, similarly to what has been discussed in Sections III.1.1 and III.1.2. These include axion-photon couplings as well as scalar or vector couplings that induce displacements or charge effects on the interferometer mirrors. In practice, because data are typically divided into segments with $T_{\text{FFT}}\sim T_{\text{coh}}$, the analytic likelihood approach can become unstable. Consequently, numerical simulations are used to estimate the sensitivity and derive robust upper limits, as done in a search on O3 KAGRA data Abac et al. (2024), which will be discussed in Section III.2.
@@ III.1.5 Distinguishing amongst dark-matter models
Only a few works Miller et al. (2022b); Manita et al. (2024) have explored ways to determine whether a particular type of DM interacted with GW detectors. In Miller et al. (2022b), the authors employ the Wiener filter Wiener et al. (1964) to compute the cross power across different detector pairs during follow-up stages, using different $T_{\text{FFT}}$ to exploit the fact that different types of DM couple differently to different detector baselines.
@@
Miller et al. (2022b) showed that the Wiener filter produces small residuals when the data contain a signal consistent with the model and large residuals when the wrong signal is applied. This property allowed the method to confidently veto spurious candidates in real O3 data. Conversely, this means that the method has the potential to confirm the detection of a certain type of DM interaction signal. This is in contrast to other methods and previous searches Guo et al. (2019); Miller et al. (2021a); Abbott et al. (2022c), which focused mostly on rejecting outliers that were shown to be compatabile with noise.
After the work of Miller et al. (2022b), Ref. Manita et al. (2024) uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different. In particular, Ref. Manita et al. (2024) how the cross-correlation SNR would change as a function of the type of DM and which baseline was considered. In both cases, cross power is important, since the individual power spectra from one detector are indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
@@ III.2 Ground- and space-based interferometers
Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons Abbott et al. (2022c); Vermeulen et al. (2021); Göttel et al. (2024); Abac et al. (2024). Furthermore, LISA Pathfinder data were also used to set constraints on dark photon DM Miller and Mendes (2023); Frerick et al. (2024). The theoretical framework for each particle is described in Section II.3 and Section II.4, respectively. In this subsection, we summarize key constraints from existing ground-based interferometers on these two particles.
@@ III.2.1 LIGO, Virgo and GEO600
In Ref. Abbott et al. (2022c), two complementary methods were employed to search for spin-1 ultralight dark matter that can couple to baryons in the test masses of LIGO in O3 data – cross-correlation and excess power, both of which are described in Section III.1.1 and Section III.1.2, respectively. Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency excess power method. On the other hand, the excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every one-Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
@@
Both searches returned no significant DM candidate signals; thus, 95% confidence-levle upper limits were set on the coupling strength $\epsilon^{2}$ of dark photons to baryons for both methods, as shown in Fig. 4.. We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments. These upper limits were derived for cross-correlation and excess power following the procedure outlined in Section III.1.1 and Section III.1.2, respectively. Both methods employed the Feldman-Cousins approach Feldman and Cousins (1998) to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller et al. (2021a).
@@
Figure 4: Adapted from Abbott et al. (2022c). Upper limits obtained from analyzing LIGO O3 data on the square of the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors. The physics behind this form of DM has been discussed in Section II.4. The limits derived from each of the methods discussed in Sections III.1.1 and III.1.2 are show in red (cross-correlation) and cyan (excess power), respectively. MICROSCOPE Bergé et al. (2018) and Eöt-Wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger et al. (2008). To produce limits on the square of the dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
@@
While cross-correlation and excess power analyses were used for dark photon searches, the LPSD technique (described in Section III.1.3) was employed in Ref. Göttel et al. (2024) to search for scalar dilaton dark matter in the LIGO O3 dataset. This analysis placed strong limits on the coupling of scalar dark matter to electrons and photons. In Fig. 5, we can see, in blue, the constraints that have been derived, which supersede, at least at low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen et al. (2021). The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search. However, at high frequencies, we can see that GEO600 still outperforms LIGO, primarily because of sophisticated quantum technologies that were employed in its design. These upper limits were derived by following the procedure outlined in Section III.1.3 and Section III.1.2.
@@
Figure 5: Taken from Göttel et al. (2024). 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors. The physics behind this form of DM has been discussed in Section II.3.1.a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively. The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen et al. (2021); Aiello et al. (2022); Aharony et al. (2021); Savalle et al. (2021); Kennedy et al. (2020); Antypas et al. (2019, 2021); Tretiak et al. (2022); Oswald et al. (2022); Campbell et al. (2021); Beloy et al. (2021); Zhang et al. (2023a); Fukusumi et al. (2023); Sherrill et al. (2023) are shown in thin grey, and constraints from searches for “fifth forces” Bergé et al. (2022); Hees et al. (2018) are depicted by the dashed red lines.

### v5

@@ II.6.1 Minimally coupled dark matter
A commonly studied example of a minimally coupled scalar DM field $\phi$ is described by the Lagrangian Khmelnitsky:2013lxt:
@@
Scalar DM can cause the arrival times of pulses from millisecond pulsars on earth to change, in an analogous way that GWs do Khmelnitsky:2013lxt, i.e. introduce stochastic fluctuations of the metric Kim:2023kyy; Kim:2023pkx. Such a difference in the arrival times can be written as Khmelnitsky:2013lxt; Porayko:2018sfa:
@@ II.6.2 Non-minimally coupled (conformal) dark matter
Another form of DM that couples only to gravity would be conformal DM. This kind of DM typically involve scalar fields that couple to the trace of the energy-momentum tensor. These models are closely related to the dilaton-type scalar DM discussed in Section II.3, but with specific couplings motivated by scalar–tensor theories of gravity, such as Fierz–Jordan–Brans–Dicke (FJBD) and Damour–Esposito–Farese (DEF) theories Fierz:1939ix; Fierz:1956zz; Dicke:1961gz; Jordan:1959eg; Damour:1992we; Damour:1993hw.
In such models, DM couples universally to the gravitational sector through a conformal rescaling of the metric. This coupling can induce a gravity-mediated force between neutron stars and the ultralight dark matter field, altering the moment of inertia of rotating pulsars – a manifestation of the Nordtvedt effect Nordtvedt:1968qs. These variations lead to deterministic changes in pulsar ToA and constitute a violation of the equivalence principle. This kind of conformal scalar DM has been extensively studied and constrained using binary pulsar timing measurements Damour:1991rd; Kramer:2006nb, which will be discussed inSection III.3.
@@
where $s_{I}$ is the angular moment sensitivity parameter computed in Kuntz:2024jxo and is of $\mathcal{O}(1)$ in FJBD theory, $\alpha$ is the strength of the coupling, $\theta(\vec{x})$ is a random phase, and $\hat{\phi}$ is a stochastic parameter whose value depends on whether the correlated, uncorrelated or pulsar correlated scenarios are considered (see Section III.3 for a description of these terms).
@@ III.1 Methods
As explained in Section II.1, the expected signal frequency is fixed by the DM mass, but has some stochastic variations of $\mathcal{O}(10^{-6})$ Centers:2019dyn. Because the interferometers always exist in the DM field, the signal is always present. Thus, we can describe the expected signal as: quasi-monochromatic and quasi-infinite duration. Essentially, we are looking for a resonance at a particular frequency fixed by the DM mass.
@@ III.1.1 Cross-correlation
Conceptually, cross-correlation Pierce:2018xmy; Guo:2019ker requires at least two separate time-domain datasets that are Fourier transformed and multiplied together to compute the cross-power in each frequency bin. The cross power is then divided by the auto-power (the power spectral density) of each detector, and then summed over all the Fourier transforms to arrive at a measure of power at each frequency. It is then divided by the standard deviation of the noise to compute the signal-to-noise ratio (SNR).
Mathematically, the cross-correlated signal strength for detector pair $IJ$ is Pierce:2018xmy; Guo:2019ker:
@@
In addition to the signal strength, we have to obtain an estimate of the variance of the noise data. This can be computed from the individual power spectral densities in each FFT as Guo:2019ker:
@@
Depending on the type of ultralight dark matter being searched for, the detectability of a signal in cross-correlation searches between interferometers depends on the spin of the DM: spin-0 (scalar), spin-1 (vector), or spin-2 (tensor) Manita:2023mnc. The SNR for a given detector pair is determined by the DM energy density fraction in that spin state, the strength of its couplings to standard-model particles, the observation time, and the relative orientation, location and geometry of the detectors. The latter is characterized by an “overlap reduction function” (ORF), which determines the sensitivity loss when applying cross-correlation to non-collocated, non-aligned detector pairs. For these models of DM, two kinds of ORFs arise: one from the spatial displacement of mirrors, and another from the finite-light time travel effects of the signal on the mirrors, a previously explained in Section II.4 Morisaki:2020gui. In particular, cross-correlation is not nearly as sensitive to the finite-light travel time effect as to the spatial effect when it was applied to search for spin-1 DM, which will be shown in Section III.2.
Larger couplings, longer effective observation times, and favorable overlap between detectors increase the SNR and thus the probability of detecting a signal. Conversely, weaker couplings, shorter observations, or misaligned detectors reduce the SNR, making a signal more difficult to distinguish from instrumental noise. After an SNR is computed via Eq. 68, it can be translated directly from SNR to coupling strengths using Eq. 120-122 in Manita:2023mnc).
@@
$h_{0,j}=\left(\frac{2\,\mathrm{SNR}}{\gamma}\right)^{1/2}\left(\frac{\,P_{I,j}\,P_{J,j}}{T_{\text{obs}}\,T_{\text{FFT}}}\right)^{1/4},$ (68)
where $\gamma$ is the ORF that depends on the way that different kinds of DM couple to the interferometers Manita:2023mnc. By inserting the SNR returned from a search (Eq. 67), we can obtain an estimate of the minimum detectable strain amplitude for a potential DM signal, and then map those amplitude values to constraints on DM couplings in spin-0, spin-1 or spin-2 using the strain amplitude equations derived in Section II, analogously to translating the SNR directly to coupling strengths in Manita:2023mnc.
III.1.2 BSD Excess power method
Cross-correlation is able to capture the phase information in each FFT; however, at the moment the software has not yet been implemented to change $T_{\text{FFT}}$ to match $T_{\text{coh}}$, which would result in optimal sensitivity to each DM mass. Thus, another method Miller:2020vsl called “BSD excess power” was developed that can vary $T_{\text{FFT}}$ as a function of frequency that employs Band-Sampled Data (BSD) structures that allow easy changes of $T_{\text{FFT}}$ piccinibsd, and that is more robust against noise disturbances. It relies on creating time-frequency spectrograms in which frequencies in each FFT are only kept if their equalized powers are (1) above a given threshold $\theta_{\rm thr}=2.5$ and (2) local maxima. The equalized power is computed at each time and frequency bin by calculating the ratio $R_{ij}$ of the square modulus of the FFT with a running median estimation of the power spectral density, as in Section III.1.1:
@@ III.1.2 BSD Excess power method
where $n$ is the number of peaks at a given frequency, and $\mu$ and $\sigma$ are the mean and standard deviation of the number of peaks in the histogram. The $CR$, like the SNR, follows a normal distribution. Using the $CR$ and assuming Gaussian noise, the minimum strain amplitude of a sinusoidal signal can be derived that would, in a frequentest interpretation, produce a detectable signal in a fraction $\geq\Gamma$ of a repeated number of experiments Astone:2014esa; Miller:2020vsl:
@@
The $\mathcal{O}(1)$ differences arise from averages over the spin-1 or spin-2 polarizations, and are valid for ground-based L-shaped GW interferometers (for triangular shapes, the averages will be slightly different). The $\frac{\sqrt{2}}{3}$ factor comes from an average over the DM Maxwell-Boltzmann distributed velocities as discussed in Section II.4, while the $\frac{2\pi}{2.4308}$ factor comes from convolving a sinusoid with a rectangular window function when computing the sensitivity of GW interferometers to a sinusoidal signal Astone:2014esa.
@@ III.1.3 Logarithmic power spectral density (LPSD) method
To maximize sensitivity to ultralight dark-matter signals, it is important to let $T_{\text{FFT}}$ match $T_{\text{coh}}$ as the DM mass varies. The logarithmic power spectral density (LPSD) method trobs2006improved; Vermeulen:2021epa; Gottel:2024cfj implements this idea by constructing spectra that are logarithmically spaced in frequency, in contrast to the BSD excess power method in Section III.1.2 that use linearly spaced spectra every 1 Hz. While this approach is computationally intensive due to the need for a separate implementation from the standard discrete FFT, it captures subtle features in the data that could otherwise be missed. In particular, since ultralight dark matter induces fractional frequency variations on the order of $\mathcal{O}(10^{-6})$, a logarithmic spectral estimate is well-suited to resolving such tiny modulations.
@@
By applying Eq. 74 to the logarithmically estimated spectra, one can extract an estimate of the strain $h(\omega)$ from the data. This estimate can then be mapped to constraints on the coupling constants $\Lambda_{\gamma}$ and $\Lambda_{e}$, providing a direct connection between the interferometer measurement and the underlying ultralight dark matter model Gottel:2024cfj; Cowan:2010js. In this way, the LPSD method offers a way to detect logarithmically-spaced frequency modulations that might be missed by linear spectral analyses.
@@ III.1.4 Stochastic summing method
The stochastic summing method Nakatsuka:2022gaf exploits the fact that when $T_{\text{FFT}}>T_{\text{coh}}$, the signal power from ultralight dark matter is spread stochastically over a range of frequencies rather than concentrated at a single frequency. Instead of searching for a narrow peak, this method sums the power across all frequency bins where the signal is expected, producing a detection statistic that captures the total available signal power.
@@
The sensitivity of the stochastic summing method has been studied for both axions and dark photons Nakatsuka:2022gaf, using a combination of analytical estimates and numerical simulations. In the absence of a signal and assuming Gaussian noise, a detection threshold can be established based on a chosen false-alarm probability. This threshold can then be used to determine, at a given confidence level, the minimum detectable signal amplitude, similar to Eq. 71 in Section III.1.2.
The exact sensitivity depends on the spectral shape of the signal, which in turn is influenced by the relative motion of the DM with respect to the interferometer. Conservative and optimistic assumptions about the direction of the DM velocity lead to $\mathcal{O}(1)$ variations in the sensitivity, which are accounted for in this method Nakatsuka:2022gaf.
Once the minimum detectable strain amplitude is determined, it can be translated into upper limits on the relevant coupling constants for each type of dark matter interaction, similarly to what has been discussed in Sections III.1.1 and III.1.2. These include axion-photon couplings as well as scalar or vector couplings that induce displacements or charge effects on the interferometer mirrors. In practice, because data are typically divided into segments with $T_{\text{FFT}}\sim T_{\text{coh}}$, the analytic likelihood approach can become unstable. Consequently, numerical simulations are used to estimate the sensitivity and derive robust upper limits, as done in a search on O3 KAGRA data KAGRA:2024ipf, which will be discussed in Section III.2.
@@ III.1.5 Distinguishing amongst dark-matter models
Only a few works Miller:2022wxu; Manita:2023mnc have explored ways to determine whether a particular type of DM interacted with GW detectors. In Miller:2022wxu, the authors employ the Wiener filter wiener1964extrapolation to compute the cross power across different detector pairs during follow-up stages, using different $T_{\text{FFT}}$ to exploit the fact that different types of DM couple differently to different detector baselines.
@@
Miller:2022wxu showed that the Wiener filter produces small residuals when the data contain a signal consistent with the model and large residuals when the wrong signal is applied. This property allowed the method to confidently veto spurious candidates in real O3 data. Conversely, this means that the method has the potential to confirm the detection of a certain type of DM interaction signal. This is in contrast to other methods and previous searches Guo:2019ker; Miller:2020vsl; LIGOScientific:2021odm, which focused mostly on rejecting outliers that were shown to be compatabile with noise.
After the work of Miller:2022wxu, Ref. Manita:2023mnc uses standard cross correlation to argue that the overlap reduction function can be used as way to distinguish amongst different spins of DM particles, since the cross power will be different. In particular, Ref. Manita:2023mnc how the cross-correlation SNR would change as a function of the type of DM and which baseline was considered. In both cases, cross power is important, since the individual power spectra from one detector are indistinguishable for all types of ultralight dark matter, which arises because the phase evolutions for the considered types of ultralight dark matter here are the same.
@@ III.2 Ground- and space-based interferometers
Searches for DM particles interacting with the LIGO, Virgo, KAGRA and GEO600 interferometers have been performed using data from the most recent observing run, O3, resulting in constraints on dilatons and dark photons LIGOScientific:2021odm; Vermeulen:2021epa; Gottel:2024cfj; KAGRA:2024ipf. Furthermore, LISA Pathfinder data were also used to set constraints on dark photon DM Miller:2023kkd; Frerick:2023xnf. The theoretical framework for each particle is described in Section II.3 and Section II.4, respectively. In this subsection, we summarize key constraints from existing ground-based interferometers on these two particles.
@@ III.2.1 LIGO, Virgo and GEO600
In Ref. LIGOScientific:2021odm, two complementary methods were employed to search for spin-1 ultralight dark matter that can couple to baryons in the test masses of LIGO in O3 data – cross-correlation and BSD excess power, both of which are described in Section III.1.1 and Section III.1.2, respectively. Cross-correlation benefits from utilizing the phase information of the signal, something that is lost when using the time-frequency BSD excess power method. On the other hand, the BSD excess power method matches its coherence time $T_{\text{FFT}}$ to the signal coherence time $T_{\text{coh}}$ in every one-Hz band analyzed, allowing an improved sensitivity across the frequency range compared to cross-corelation, which fixed $T_{\text{FFT}}=1800$ s.
@@
Both searches returned no significant DM candidate signals; thus, 95% confidence-level upper limits were set on the coupling strength $\epsilon^{2}$ of dark photons to baryons for both methods, as shown in Fig. 4.. We can see, across a broad frequency range, that results from GW interferometers surpass by orders of magnitude constraints from Eöt-Wash and the MICROSCOPE experiments. These upper limits were derived for cross-correlation and BSD excess power following the procedure outlined in Section III.1.1 and Section III.1.2, respectively. Both methods employed the Feldman-Cousins approach Feldman:1997qc to set upper limits, which is robust against non-Gaussian noise and has been shown to produce conservative upper limits with respect to those that would have been obtained through simulations Miller:2020vsl.
@@
Figure 4: Adapted from LIGOScientific:2021odm. Upper limits obtained from analyzing LIGO O3 data on the square of the coupling of dark photons to baryons $U(1)_{\rm B}$ in the LIGO mirrors. The physics behind this form of DM has been discussed in Section II.4. The limits derived from each of the methods discussed in Sections III.1.1 and III.1.2 are show in red (cross-correlation) and cyan (BSD excess power), respectively. MICROSCOPE Berge:2017ovy and Eöt-Wash torsion balance upper limits are plotted as a comparison to the results here Schlamminger:2007ht. To produce limits on the square of the dark photon/baryon-lepton coupling, $U(1)_{\rm B-L}$ in the LIGO mirrors, these limits should be multiplied by four.
@@
While cross-correlation and BSD excess power analyses were used for dark photon searches, the LPSD technique (described in Section III.1.3) was employed in Ref. Gottel:2024cfj to search for scalar dilaton dark matter in the LIGO O3 dataset. This analysis placed strong limits on the coupling of scalar dark matter to electrons and photons. In Fig. 5, we can see, in blue, the constraints that have been derived, which supersede, at least at low masses, those that were derived in a GEO600 search for the same kind of DM Vermeulen:2021epa. The enhancement comes from the improved low-frequency sensitivity of LIGO relative to GEO600, as well as incorporating the interaction of scalar DM with the mirrors, as well as the beam splitter, into the search. However, at high frequencies, we can see that GEO600 still outperforms LIGO, primarily because of sophisticated quantum technologies that were employed in its design. These upper limits were derived by following the procedure outlined in Section III.1.3 and Section III.1.2.
@@
Figure 5: Taken from Gottel:2024cfj. 95% confidence-level upper limits on $\Lambda_{i}^{-1}$, the coupling of dilaton DM to electrons or photons, as a function of mass and frequency from LIGO O3 data, which would have cause time-dependent oscillations of the sizes and indices of refraction of the beamsplitter and the LIGO mirrors. The physics behind this form of DM has been discussed in Section II.3.1.a) and b) show these constraints compared to existing ones on $\Lambda_{e}$ and $\Lambda_{\gamma}$, respectively. The results from the LIGO O3 search are shown by the thick blue line, constraints from direct experimental searches for DM Vermeulen:2021epa; Aiello:2021wlp; Aharony:2019iad; Savalle:2020vgz; Kennedy:2020bac; Antypas:2019qji; Antypas:2020rtg; Tretiak:2022ndx; Oswald:2021vtc; Campbell:2020fvq; BACON:2020ubh; Zhang:2022ewz; Fukusumi:2023kqd; Sherrill:2023zah are shown in thin grey, and constraints from searches for “fifth forces” Berge:2021yye; Hees:2018fpg are depicted by the dashed red lines.